# Neural Topic Modeling with Large Language Models in the Loop

### v2

@@ Abstract
Topic modeling is a fundamental task in natural language processing, allowing the discovery of latent thematic structures in text corpora. While Large Language Models (LLMs) have demonstrated promising capabilities in topic discovery, their direct application to topic modeling suffers from issues such as incomplete topic coverage, misalignment of topics, and inefficiency. To address these limitations, we propose LLM-ITL, a novel LLM-in-the-loop framework that integrates LLMs with Neural Topic Models (NTMs). In LLM-ITL, global topics and document representations are learned through the NTM. Meanwhile, an LLM refines these topics using an Optimal Transport (OT)-based alignment objective, where the refinement is dynamically adjusted based on the LLM’s confidence in suggesting topical words for each set of input words. With the flexibility of being integrated into many existing NTMs, the proposed approach enhances the interpretability of topics while preserving the efficiency of NTMs in learning topics and document representations. Extensive experiments demonstrate that LLM-ITL helps NTMs significantly improve their topic interpretability while maintaining the quality of document representation. Our code and datasets will be available at Github.
@@
Xiaohao Yang1, He Zhao2, Weijie Xu3, Yuanyuan Qi1, Jueqing Lu1, Dinh Phung1, Lan Du1 1Faculty of IT, Monash University, Australia 2CSIRO’s Data61, Australia  3Amazon, America {xiaohao.yang,yuanyuan.qi,jueqing.lu,dinh.phung,lan.du}@monash.edu he.zhao@data61.csiro.au,  weijie.xu@amazon.com
@@ 1 Introduction
Topic modeling is an essential task in natural language processing that uncovers hidden thematic structures within text collections in an unsupervised way. The ability to automatically extract topics has proven to be invaluable across a range of disciplines, such as bioinformatics (Liu et al., 2016), marketing research (Reisenbichler and Reutterer, 2019), and information retrieval (Yi and Allan, 2009). Topic models are conventionally based on probabilistic frameworks that generate a set of interpretable global topics, each represented as a distribution over vocabulary terms as Latent Dirichlet Allocation (LDA) (Blei et al., 2003). These topics are then used to represent individual documents as mixtures of topics, providing a structured and interpretable view of the corpus. Recently, research on Neural Topic Models (NTMs) (Zhao et al., 2021; Churchill and Singh, 2022; Abdelrazek et al., 2023; Wu et al., 2024a) has been popular, which uses deep neural networks to model document-topic distributions, enabling more expressive and flexible representations compared to their probabilistic counterparts.
While Large Language Models (LLMs) (OpenAI, 2022; Touvron et al., 2023a, b) have redefined the landscape of natural language processing, topic models can still be valuable tools for text analysis. Specifically, LLMs can provide a fine-grained understanding of a document; however, given a large collection of domain-specific documents, topic models are more suitable to obtain a clear global view of the topics in a more interpretable way with much less computational cost. Unsurprisingly, it has been a trending research direction to use LLMs to improve topic modeling (Rijcken et al., 2023; Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024; Chang et al., 2024). Despite the promising performance of these initial studies, most existing methods involve prompting LLMs to generate topics for each document in the corpus, which may lead to several limitations. As LLMs are asked to focus on a document individually, they may be unable to cover global topics across all the documents in the corpus (Doi et al., 2024), which is critical in topic modeling. Moreover, although LLMs excel at capturing local context, they usually struggle with long documents with multiple interrelated topics, which may evolve or shift throughout the text. With their limited window of focus, LLMs may miss key topics of a document that are necessary to fully understand its content. Finally, it is computationally expensive as LLMs have to do inference for every document in the corpus or subsets; thus, existing methods usually scale poorly with large datasets.
@@ 2.2 Neural Topic Models
A Neural Topic Model (NTM) is typically trained by modeling $p(\bm{z}|\bm{x})$ and $p(\bm{x}|\bm{z})$, where $\bm{x}\in\mathbb{N}^{V}$ represents the Bag-of-Words (BOWs) of a document. NTMs, which employ deep neural networks for topic modeling, are commonly based on Variational Auto-Encoders (VAEs) (Kingma and Welling, 2013) and Amortized Variational Inference (AVI) (Rezende et al., 2014). For VAE-NTMs, $p(\bm{x}|\bm{z})$ is modeled by a decoder network $\bm{\phi}$, i.e., $\bm{x}:=\bm{\phi}(\bm{z})$. The posterior $p(\bm{z}|\bm{x})$ is approximated by $q(\bm{z}|\bm{x})$, which is modeled by an encoder network $\bm{\theta}$, i.e., $\bm{z}:=\bm{\theta}(\bm{x})$. The training objective of VAE-NTMs is to maximize the Evidence Lower Bound (ELBO):
@@ 2.3 Optimal Transport
Optimal Transport (OT) has been widely used for comparing probability distributions (Cuturi, 2013; Frogner et al., 2015; Seguy et al., 2017; Peyré et al., 2019). Let $\mu(\bm{x},\bm{a}):=\sum_{i=1}^{N}a_{i}\delta_{x_{i}}$ and $\mu(\bm{y},\bm{b}):=\sum_{j=1}^{M}b_{j}\delta_{y_{j}}$ be two discrete distributions, where $\bm{a}:=[a_{1},\ldots,a_{N}]$ and $\bm{b}:=[b_{1},\ldots,b_{M}]$ are the probability vectors; $\bm{x}:=\{x_{1},\ldots,x_{N}\}$ and $\bm{y}:=\{y_{1},\ldots,y_{M}\}$ are the supports of these two distributions. The OT distance between $\mu(\bm{x},\bm{a})$ and $\mu(\bm{y},\bm{b})$ is obtained by finding the optimal transport plan $\bm{P}^{*}$ that transports the probability mass from $\bm{a}\in\Delta^{N}$ to $\bm{b}\in\Delta^{M}$, written as following:
@@ 3.1 LLM-based Topic Suggestion
where $\bm{w}$ represents the original topic words; $\bm{\theta}^{\text{llm}}$ denotes the LLM model; the topic label $\bm{w}^{l}$ and the refined words $\bm{w}^{\prime}$ are extracted as subsequences from the LLM’s output $\bm{s}$. Notably, the words in $\bm{w}^{\prime}$ are further filtered based on the corpus vocabulary to ensure that no out-of-vocabulary (OOV) words are included. Our used prompt is illustrated in Figure A1. A study of prompt variants is provided in Appendix H.
@@ 3.2 OT-based Topic Alignment
A key innovation in LLM-ITL is the use of Optimal Transport (OT) distance to align the topic word distributions generated by the NTM with the refined topic word distributions provided by the LLM. OT is a mathematical framework that computes the “cost” of transforming one probability distribution into another, and has shown its effectiveness in measuring the alignment between two sets of words (Kusner et al., 2015; Yang et al., 2024).
@@ 4 Related Work
Classical topic models, such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and its variants (Blei and Lafferty, 2006; Rosen-Zvi et al., 2012; Yan et al., 2013), are Bayesian probabilistic models with various generative assumptions about the documents. Neural Topic Models (NTMs) (Miao et al., 2017; Srivastava and Sutton, 2017; Card et al., 2017; Dieng et al., 2020; Bianchi et al., 2020; Zhao et al., 2020; Nguyen and Luu, 2021; Shen et al., 2021; Duan et al., 2021; Li et al., 2022; Xu et al., 2023a, b; Yang et al., 2023; Miyamoto et al., 2023; Wu et al., 2024b; Chen et al., 2025a, b) use deep neural networks to learn topics and document representations. Clustering-based topic models (Sia et al., 2020; Grootendorst, 2022; Angelov and Inkpen, 2024) discover topics using clustering algorithms based on embeddings from pre-trained language models. Ours is not a specific topic model, but rather a general framework designed to integrate LLMs with a wide range of NTMs.
@@
LLMs have been involved in topic modeling in various ways. Rijcken et al. (2023) investigate the use of ChatGPT (OpenAI, 2022) to generate descriptions for topic words and found the effectiveness of these topic descriptions. Recent works leverage LLMs for topic model evaluation in different ways, such as applying LLMs for word intrusion or topic rating for topics (Rahimi et al., 2023; Stammbach et al., 2023), or keyword generation for documents (Yang et al., 2024). LLM-based topic models have emerged (Wang et al., 2023; Pham et al., 2023; Mu et al., 2024; Doi et al., 2024), which prompt LLMs to generate topics and assign topics to documents. Unlike existing methods that focus on document-level analysis relying on LLMs, our approach prompts LLMs to analyze a group of topic words, which are then used to guide the training of NTMs. More recently, Chang et al. (2024) show that LLMs are effective at refining topic words, leading to improved topic coherence. However, their method refines the topic words of trained topic models in a post-hoc manner, whereas ours incorporates refinement as a regularization term during training. Our method is also loosely related to uncertainty estimation of LLMs, and we omit the discussion on this in Appendix C.
@@ 5.1 Experimental Setup
We conduct experiments on four widely used and publicly available datasets for topic modeling, including 20Newsgroup (Lang, 1995) (20News), Reuters-21578 (Aletras and Stevenson, 2013) (R8), DBpedia (Auer et al., 2007) and AGNews (Zhang et al., 2015). Further details of these datasets are described in the Appendix D.1. The number of mined topics (i.e., $K$) is commonly regarded as a hyper-parameter for the dataset (Zhao et al., 2020; Wu et al., 2024a). For datasets containing long documents, such as 20News and R8, we set the number of topics to 50. For datasets with short documents, such as DBpedia and AGNews, we set the number to 25. We also run experiments at different $K$ values, which are reported in Appendix E.
@@
The LLM-ITL framework is highly modular and can be seamlessly integrated with a wide range of NTMs. We integrated LLM-ITL with 8 commonly benchmarked NTMs, as listed in Table 1. We also evaluate other types of topic models, including probabilistic models such as LDA (Blei et al., 2003), cluster-based models like BERTopic (Grootendorst, 2022), and the latest LLM-based model, TopicGPT (Pham et al., 2023). Further details about these baselines and their settings are provided in Appendix D.2.
@@
Evaluation Metrics
We evaluate both the topic quality and the document representation quality for topic models. For topic quality, we apply the widely used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015). We report the average $C_{V}$ values of all learned topics. For document representation quality, we evaluate the alignment between a document’s true label and the top-weighted topic of its topical representation, known as topic alignment (Chuang et al., 2013; Pham et al., 2023). This is commonly evaluated using external clustering metrcis, Purity and Normalized Mutual Information (NMI). Since Purity and NMI are considered equally important, fall within the same range (from 0 to 1), and are often reported together, we report their average as PN, serving as an overall indicator of topic alignment performance. Detailed results for Purity and NMI are provided in Appendix E.5. Intuitively, topic coherence ($C_{V}$) reflects how coherent the learned topic words are, while topic alignment (PN) indicates how well the model represents the documents through the learned topics. Further details on the calculation of these metrics are provided in Appendix D.3. In addition, we compute the performance improvement of LLM-ITL over its base model as: $\frac{S(\text{base+LLM-ITL})-S(\text{base})}{S(\text{base})}$, where $S(\cdot)$ represents the evaluation metric. As for other topic model evaluation metrics, including topic diversity (TD) (Dieng et al., 2020) and overall topic quality (TQ) (Wang et al., 2022) are reported in Appendix E.3 and E.4, respectively.
5.2 Results
@@
Topic Coherence and Alignment
We show the performance of topic coherence and topic alignment for different models in Table 1. We summarize the following remarks based on the results: (1) For topic coherence, using LLM-ITL significantly improves performance across all cases, with minimum gains of +7.4% (NSTM as the base model on DBpedia) and maximum gains of +70.6% (ECRTM as the base model on 20News) over the base model. (2) In terms of topic alignment, integrating LLM-ITL demonstrates comparable performance to its base model, with changes ranging from -3.5% to +4.2%. (3) Moreover, for long-document corpora such as 20News and R8, applying LLM-ITL outperforms existing LLM-based topic models like TopicGPT in terms of topic alignment in most cases. This suggests that LLMs alone struggle to fully capture the topics of long documents, highlighting the advantages of integrating NTMs to enhance topical representation for document collections. Additional results with different settings and evaluation metrics are presented in Appendix E.
Balancing Topic Coherence and Alignment

### v3

@@ Abstract
Topic modeling is a fundamental task in natural language processing, allowing the discovery of latent thematic structures in text corpora. While Large Language Models (LLMs) have demonstrated promising capabilities in topic discovery, their direct application to topic modeling suffers from issues such as incomplete topic coverage, misalignment of topics, and inefficiency. To address these limitations, we propose LLM-ITL, a novel LLM-in-the-loop framework that integrates LLMs with Neural Topic Models (NTMs). In LLM-ITL, global topics and document representations are learned through the NTM. Meanwhile, an LLM refines these topics using an Optimal Transport (OT)-based alignment objective, where the refinement is dynamically adjusted based on the LLM’s confidence in suggesting topical words for each set of input words. With the flexibility of being integrated into many existing NTMs, the proposed approach enhances the interpretability of topics while preserving the efficiency of NTMs in learning topics and document representations. Extensive experiments demonstrate that LLM-ITL helps NTMs significantly improve their topic interpretability while maintaining the quality of document representation. Our code and datasets are available at https://github.com/Xiaohao-Yang/LLM-ITL
@@
Xiaohao Yang1, He Zhao2thanks: Corresponding authors: Lan Du, He Zhao, Weijie Xu3, Yuanyuan Qi1, Jueqing Lu1, Dinh Phung1, Lan Du1 1Faculty of IT, Monash University, Australia 2CSIRO’s Data61, Australia  3Amazon, America {xiaohao.yang,yuanyuan.qi,jueqing.lu,dinh.phung,lan.du}@monash.edu he.zhao@data61.csiro.au,  weijie.xu@amazon.com
@@ 1 Introduction
Topic modeling is an essential task in natural language processing that uncovers hidden thematic structures within text collections in an unsupervised way. The ability to automatically extract topics has proven to be invaluable across a range of disciplines, such as bioinformatics (Liu et al., 2016), marketing research (Reisenbichler and Reutterer, 2019), and information retrieval (Yi and Allan, 2009). Topic models are conventionally based on probabilistic frameworks such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and its hierarchical Bayesian extensions (Paisley et al., 2015; Gan et al., 2015; Zhou et al., 2016; Zhao et al., 2018b, a), which generate a set of interpretable global topics, each represented as a distribution over vocabulary terms. These topics are then used to represent individual documents as mixtures of topics, providing a structured and interpretable view of the corpus. Recently, research on Neural Topic Models (NTMs) (Zhao et al., 2021a; Churchill and Singh, 2022; Abdelrazek et al., 2023; Wu et al., 2024a) has been popular, which uses deep neural networks to model document-topic distributions, enabling more expressive and flexible representations compared to their probabilistic counterparts.
While Large Language Models (LLMs) (OpenAI, 2023; Touvron et al., 2023a, b) have redefined the landscape of natural language processing, topic models can still be valuable tools for text analysis. Specifically, LLMs can provide a fine-grained understanding of a document; however, given a large collection of domain-specific documents, topic models are more suitable to obtain a clear global view of the topics in a more interpretable way with much less computational cost. Unsurprisingly, it has been a trending research direction to use LLMs to improve topic modeling (Rijcken et al., 2023; Wang et al., 2023; Pham et al., 2024; Mu et al., 2024; Doi et al., 2024; Chang et al., 2024). Despite the promising performance of these initial studies, most existing methods involve prompting LLMs to generate topics for each document in the corpus, which may lead to several limitations. As LLMs are asked to focus on a document individually, they may be unable to cover global topics across all the documents in the corpus (Doi et al., 2024), which is critical in topic modeling. Moreover, although LLMs excel at capturing local context, they usually struggle with long documents with multiple interrelated topics, which may evolve or shift throughout the text. With their limited window of focus, LLMs may miss key topics of a document that are necessary to fully understand its content. Finally, it is computationally expensive as LLMs have to do inference for every document in the corpus or subsets; thus, existing methods usually scale poorly with large datasets.
@@ 2.2 Neural Topic Models
A Neural Topic Model (NTM) is typically trained by modeling $p(\bm{z}|\bm{x})$ and $p(\bm{x}|\bm{z})$, where $\bm{x}\in\mathbb{N}^{V}$ represents the Bag-of-Words (BoWs) of a document. NTMs, which employ deep neural networks for topic modeling, are commonly based on Variational Auto-Encoders (VAEs) (Kingma and Welling, 2014) and Amortized Variational Inference (AVI) (Rezende et al., 2014). For VAE-NTMs, $p(\bm{x}|\bm{z})$ is modeled by a decoder network $\bm{\phi}$, i.e., $\bm{x}:=\bm{\phi}(\bm{z})$. The posterior $p(\bm{z}|\bm{x})$ is approximated by $q(\bm{z}|\bm{x})$, which is modeled by an encoder network $\bm{\theta}$, i.e., $\bm{z}:=\bm{\theta}(\bm{x})$. The training objective of VAE-NTMs is to maximize the Evidence Lower Bound (ELBO):
@@ 2.3 Optimal Transport
Optimal Transport (OT) has been widely used for comparing probability distributions (Cuturi, 2013; Frogner et al., 2015; Seguy et al., 2018; Peyré et al., 2019), which extensive applications in machine learning and related areas Ge et al. (2021); Nguyen et al. (2021); Wang et al. (2022); Guo et al. (2022); Bui et al. (2022); Vuong et al. (2023); Zhao et al. (2023); Ye et al. (2024); Vo et al. (2024); Gao et al. (2024a); Ren et al. (2024). Let $\mu(\bm{x},\bm{a}):=\sum_{i=1}^{N}a_{i}\delta_{x_{i}}$ and $\mu(\bm{y},\bm{b}):=\sum_{j=1}^{M}b_{j}\delta_{y_{j}}$ be two discrete distributions, where $\bm{a}:=[a_{1},\ldots,a_{N}]$ and $\bm{b}:=[b_{1},\ldots,b_{M}]$ are the probability vectors; $\bm{x}:=\{x_{1},\ldots,x_{N}\}$ and $\bm{y}:=\{y_{1},\ldots,y_{M}\}$ are the supports of these two distributions. The OT distance between $\mu(\bm{x},\bm{a})$ and $\mu(\bm{y},\bm{b})$ is obtained by finding the optimal transport plan $\bm{P}^{*}$ that transports the probability mass from $\bm{a}\in\Delta^{N}$ to $\bm{b}\in\Delta^{M}$, written as following:
@@ 3 Method
Figure 2: Prompt and output of topic suggestion with CoT. We take the product of token probabilities of topic label (e.g., words in red color) as the Label Token Probability. We take the proportion of intruders (e.g., words in green color) as the Word Intrusion Confidence. $N=2$ and $M=10$ in this example.
@@ 3.1 LLM-based Topic Suggestion
where $\bm{w}$ represents the original topic words; $\bm{\theta}^{\text{llm}}$ denotes the LLM model; the topic label $\bm{w}^{l}$ and the refined words $\bm{w}^{\prime}$ are extracted as subsequences from the LLM’s output $\bm{s}$. Notably, the words in $\bm{w}^{\prime}$ are further filtered based on the corpus vocabulary to ensure that no out-of-vocabulary (OOV) words are included. Our used prompt is illustrated in Figure 2. A study of prompt variants is provided in Appendix H.
@@ 3.2 OT-based Topic Alignment
A key innovation in LLM-ITL is the use of Optimal Transport (OT) distance to align the topic word distributions generated by the NTM with the refined topic word distributions provided by the LLM. OT is a mathematical framework that computes the “cost” of transforming one probability distribution into another, and has shown its effectiveness in measuring the alignment between two sets of words (Kusner et al., 2015; Yang et al., 2025).
@@ 4 Related Work
Classical topic models, such as Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and its variants (Blei and Lafferty, 2006; Rosen-Zvi et al., 2004; Yan et al., 2013), are Bayesian probabilistic models with various generative assumptions about the documents. Neural Topic Models (NTMs) (Miao et al., 2017; Srivastava and Sutton, 2017; Card et al., 2018; Dieng et al., 2020; Bianchi et al., 2021; Zhao et al., 2021b; Nguyen and Tuan, 2021; Shen et al., 2021; Duan et al., 2021; Li et al., 2022; Xu et al., 2023a, b; Yang et al., 2023; Miyamoto et al., 2023; Wu et al., 2024b; Chen et al., 2025a, b) use deep neural networks to learn topics and document representations. Clustering-based topic models (Sia et al., 2020; Grootendorst, 2022; Angelov and Inkpen, 2024) discover topics using clustering algorithms based on embeddings from pre-trained language models. Ours is not a specific topic model, but rather a general framework designed to integrate LLMs with a wide range of NTMs.
@@
LLMs have been involved in topic modeling in various ways. Rijcken et al. (2023) investigate the use of ChatGPT (OpenAI, 2023) to generate descriptions for topic words and found the effectiveness of these topic descriptions. Recent works leverage LLMs for topic model evaluation in different ways, such as applying LLMs for word intrusion or topic rating for topics (Rahimi et al., 2024; Stammbach et al., 2023), or keyword generation for documents (Yang et al., 2025). LLM-based topic models have emerged (Wang et al., 2023; Pham et al., 2024; Mu et al., 2024; Doi et al., 2024), which prompt LLMs to generate topics and assign topics to documents. Unlike existing methods that focus on document-level analysis relying on LLMs, our approach prompts LLMs to analyze a group of topic words, which are then used to guide the training of NTMs. More recently, Chang et al. (2024) show that LLMs are effective at refining topic words, leading to improved topic coherence. However, their method refines the topic words of trained topic models in a post-hoc manner, whereas ours incorporates refinement as a regularization term during training. Our method is also loosely related to uncertainty estimation of LLMs, and we omit the discussion on this in Appendix C.
@@ 5.1 Experimental Setup
We conduct experiments on four widely used and publicly available datasets for topic modeling, including 20Newsgroup (Lang, 1995) (20News), Reuters-21578 (Aletras and Stevenson, 2013) (R8), DBpedia (Auer et al., 2007) and AGNews (Zhang et al., 2015). Further details of these datasets are described in the Appendix D.1. The number of mined topics (i.e., $K$) is commonly regarded as a hyper-parameter for the dataset (Zhao et al., 2021b; Wu et al., 2024a). For datasets containing long documents, such as 20News and R8, we set the number of topics to 50. For datasets with short documents, such as DBpedia and AGNews, we set the number to 25. We also run experiments at different $K$ values, which are reported in Appendix E.
@@
The LLM-ITL framework is highly modular and can be seamlessly integrated with a wide range of NTMs. We integrated LLM-ITL with 8 commonly-used NTMs, as listed in Table 1. We also evaluate other types of topic models, including probabilistic models such as LDA (Blei et al., 2003), cluster-based models like BERTopic (Grootendorst, 2022), and the latest LLM-based model, TopicGPT (Pham et al., 2024). Further details about these baselines and their settings are provided in Appendix D.2.
@@
Evaluation Metrics
We evaluate both the topic quality and the document representation quality for topic models. For topic quality, we apply the widely used topic coherence metric, $\bm{C_{V}}$ (Röder et al., 2015). We report the average $C_{V}$ values of all learned topics. For document representation quality, we evaluate the alignment between a document’s true label and the top-weighted topic of its topical representation, known as topic alignment (Chuang et al., 2013; Pham et al., 2024). This is commonly evaluated using external clustering metrics, Purity and Normalized Mutual Information (NMI). Since Purity and NMI are considered equally important, fall within the same range (from 0 to 1), and are often reported together, we report their average as PN, serving as an overall indicator of topic alignment performance. Detailed results for Purity and NMI are provided in Appendix E.5. Intuitively, topic coherence ($C_{V}$) reflects how coherent the learned topic words are, while topic alignment (PN) indicates how well the model represents the documents through the learned topics. Further details on the calculation of these metrics are provided in Appendix D.3. In addition, we compute the performance improvement of LLM-ITL over its base model as: $\frac{S(\text{base+LLM-ITL})-S(\text{base})}{S(\text{base})}$, where $S(\cdot)$ represents the evaluation metric. As for other topic model evaluation metrics, including topic diversity (TD) (Dieng et al., 2020) and overall topic quality (TQ) (dongsheng wang et al., 2022) are reported in Appendix E.3 and E.4, respectively.
@@
5.2 Results
Topic Coherence and Alignment
We show the performance of topic coherence and topic alignment for different models in Table 1. We summarize the following remarks based on the results: (1) For topic coherence, using LLM-ITL significantly improves performance across all cases, with minimum gains of +7.4% (NSTM as the base model on DBpedia) and maximum gains of +70.6% (ECRTM as the base model on 20News) over the base model. (2) In terms of topic alignment, integrating LLM-ITL demonstrates comparable performance to its base model, with changes ranging from -3.5% to +4.2%. (3) Moreover, for long-document corpora such as 20News and R8, applying LLM-ITL outperforms existing LLM-based topic models like TopicGPT in terms of topic alignment in most cases. This suggests that LLMs alone struggle to fully capture the topics of long documents, highlighting the advantages of integrating NTMs to enhance topical representation for document collections. Additional results with different settings and evaluation metrics are presented in Appendix E and Appendix I.